# Scaling of energy delivered through an electrostatic discharge to a small series load

### v1

@@ 4.1 Gap length dependence of breakdown voltage
Since the primary objective of this project is to understand what fraction of the stored energy is delivered to a victim load, we made measurements of the breakdown voltage as a function of the electrode gap for all of the discharge conditions. Figure 7(a) shows the measured voltage just before the start of the discharge as a function of gap separation for the SVC system that had external capacitance values of $C_{x}=$ 100, 200 and 700 pF (shown in shades of blue) and 1.27 cm diameter brass spherical electrodes, as well as the OAS system where the capacitance was 700 pF (orange circles) and 3.75 cm diameter graphite spherical electrodes. The trends are close to linear in $h$ as expected: Figure 7(b) shows the breakdown field calculated by fitting the $h<2$ mm section of the data to a line. In fitting the data, we find voltage offsets of approximately $V_{0}=1.45$ kV at $h=0$. Similar offsets have previously been observed in the literature [28, 29] and likely result from a dielectric layer on the electrodes, such as an oxide layer [30]. The breakdown fields range from 28-34 kV/cm. The breakdown voltage is expected to vary with electrode geometry and material but can also depend on ambient air pressure and humidity, which was not controlled in our experiment.
@@
Figure 7: Demonstrating the effect of varying capacitance on (a) breakdown voltage, and (b) breakdown field strength for the SVC with $C_{x}$ = 100, 200, and 700 pF, $R_{v}=0.25$ $\Omega$, and brass 1.27 cm diameter electrodes. The OAS data ($C_{x}$ = 700 pF, $R_{v}=0.00514$ $\Omega$, 3.75 cm graphite sphere electrodes) is overlaid for comparison with the SVC 700 pF trace to observe the effects of circuit inductance. All data points are averaged over 10 discharge events per gap length and normalized to the victim load size of 0.1 $\Omega$. The legend in (b) applies to both plots.
@@
In Figure 8, we show similar data for the breakdown voltage vs gap length for various electrode geometries over a broader range of gap lengths. The SVC system was used with $C_{x}=100$ pF and 0.10 $\Omega$ CVR. The legend for the different electrode geometries follows the notation anode-cathode; for example, “needle-plane” indicates a needle-shaped anode and a planar cathode.
@@
Figure 8: Demonstrating the effect of varying electrode geometries on (a) breakdown voltage, and (b) breakdown field strength for the SVC with $C_{x}=100$ pF and $R_{v}=0.1$ $\Omega$. All data points are averaged over 10 discharge events per gap length. Similar $V_{0}$ offsets were found with a fit on the $V_{\text{th}}$ data with a range between 1.3 and 1.8 kV. The legend in (b) detailing anode-cathode electrode pairings applies to both plots.
@@
For the non-needle electrodes, there is a linear increase in breakdown voltage of approximately 26 kV/cm at smaller gap lengths. The breakdown voltage is lower than the linear trend for larger gap distances. We interpret this as due to the fact that for large gap lengths, the field is less uniform and concentrated near the electrodes. The needle-shaped electrode, either as the anode or cathode, emphasizes this effect, significantly reducing the breakdown voltage; this effect reduces the necessary potentials to initiate discharge events in gaps with strongly non-uniform electric field distributions [31].
@@ 4.2 Varying circuit capacitance
Figure 9: Demonstrating the effect of varying capacitance on (a) energy dissipated by $R_{v}$, (b) $\overline{\eta}_{v}$, and (c) $a_{R}$ for the SVC with $C_{x}$ = 100, 200, and 700 pF, $R_{v}=0.25$ $\Omega$, and 1.27 cm diameter brass sphere electrodes. The OAS data ($C_{x}$ = 700 pF, CVR = 0.00514 $\Omega$, 3.75 graphite sphere electrodes) is overlaid for comparison with the SVC 700 pF trace to observe the effects of circuit inductance. All data points are averaged over 10 discharge events per gap length and normalized to the victim load size of 0.1 $\Omega$. The legend in (b) applies to all three plots.
For a given gap separation, an increase in capacitance increases the initial stored energy and charge. The energy transferred to the victim load $\mathcal{E}_{v}$ is shown in Figure 9(a). As expected, $\mathcal{E}_{v}$ increases with $h$ and $C_{x}$. As detailed in Section 3, much of the variation shown can be accounted for in the parameter $\overline{\eta}_{v}$ that normalizes $\mathcal{E}_{v}$ to the $\mathcal{E}_{0}$, $R_{V}$ and $C_{x}$ (see Eq. (12)). Figure 9(b), shows that for $h>1$ mm, $\overline{\eta}_{v}\approx 0.4\ \Omega^{-1}$nF-1 for all of the tested capacitances. The fraction of stored energy transferred to the victim load is quite small: even for $C=700$ pF, $\eta_{v}$ is only $2.8\%$. This confirms the utility of the Rompe-Weizel model. There is, however, a rise in $\overline{\eta}_{v}$ at lower gaps for the lower stored energy cases. This deviation from a constant results from the voltage offset at $h=0$ giving a stored energy that is larger than expected at small gaps.
The rise in $\overline{\eta}_{v}$ with smaller gaps seen in Figure 9(b) is related to the observed breakdown voltage offset, $V_{0}$: the stored energy has a component $\mathcal{E}_{0}=\frac{1}{2}CV_{0}^{2}$ that is independent of $h$. When we account for this voltage offset, we can use a different approach to plotting the data. Eq.(11) can be adjusted to replace $E_{\text{th}}$ with $(V_{0}+E_{\text{th}}h)/h=V_{\text{th}}/h$, where $V_{\text{th}}$ is the observed breakdown threshold voltage. Making use of this relation, we can instead plot a calculated $a_{R}(h)$:
$a_{R}=\frac{4\,\eta_{V}\,h^{2}}{R_{v}C_{x}V_{\text{th}}^{2}}.$ (13)
Figure 9(c) shows this calculated value of $a_{R}(h)$. The values are fairly constant at larger gap lengths, with some slight decrease at smaller gaps. The $a_{R}$ values are in a range similar to those observed by Jobava et al. [27] that are indicated in the shaded region. We believe that the decrease in $a_{R}$ for small $h$ in these plots is related to a small additional passive resistance in series with the CVR. We will address this issue in a later paper.
@@ 4.3 Varying circuit inductance
The influence of circuit inductance on discharge dynamics was investigated by operating the SVC (124 nH, 1.27 cm diameter brass spherical electrodes) and OAS (1.1 $\mu$H, 3.75 cm diameter graphite spherical electrodes) systems with the same 700 pF external capacitor. The data for the OAS system is shown in orange circles in Figure 7. The breakdown voltage is seen to be similar, with a slightly higher offset voltage, likely resulting from the different electrode materials. Furthermore, the energy dissipated by the victim load was comparable between the SVC and OAS configurations, Figure 9(a).
Figure 9(b) displays the normalized energy transfer to the external capacitor and victim load, calculated using Eq. (12). For both the 700 pF SVC and OAS datasets, the normalized energy remained relatively constant across the gap lengths tested. This agreement in discharge dynamics between the two systems indicates that a tenfold increase in inductance does not significantly impact the energy delivered to the victim load. Since most of the stored energy is dissipated into the spark in the early phase of the discharge, the spark resistance during the latter phase of the discharge is constant. In this case, where the dynamics are in a linear regime, $\eta_{v}$ would be expected to be independent of the inductance.
@@ 4.4 Varying electrode geometry
The spherical-electrode geometry for small gaps leads to relatively uniform fields between the electrodes. When we increase the gap length and vary the electrode geometry, the shape and finite size of the electrodes result in non-uniformity of the field, as we saw in the breakdown voltages in Figure 8. The consequence of these effects can be seen in Figure 10(a), which shows the energy delivered to the victim load as a function of gap length for several combinations of electrodes and polarities. The electrodes featuring a needle-shaped electrode and the orientation plane-sphere have lower absolute energy transfer and peak current to the victim. In the plane-sphere geometry, we see lower energy transfer when the electrode is spherical: while this difference is not currently fully understood, it should be noted that in this case the electrodes are made of different materials.
Figure 10: (a) Energy dissipated by $R_{v}$, (b) $\overline{\eta}_{v}$, and (c) $a_{R}$ for the SVC with $C_{x}=100$ pF and $R_{v}=0.1$ $\Omega$, demonstrating the effect of varying electrode geometries. All data points are averaged over 10 discharge events per gap length. The legend in (b) detailing anode-cathode electrode pairings applies to all plots. In (c), the data in blue triangles involving a needle electrode have been divided by a factor of 2.
Figure 10(b) is the counterpart to Figure 9(b), showing how the fraction of stored energy transferred to the victim load scaled by $R_{v}$ and $C$ varies with gap length. The figure clearly shows that even here, the $\overline{\eta}_{v}$ scaling results in a reasonably constant value for $h>2$ mm, in spite of the different discharge dynamics. Interestingly, while the RW model predicts $\overline{\eta}_{v}$ to be a constant with no gap length dependence, each geometry shows an initial steep decrease for gaps below 1 mm, followed by a more gradually down-sloping linear relationship at longer gap lengths. The variability in the curve, which differs from a straight line, indicates that more interesting physics is present than the RW model represents, especially at the smaller gap lengths.
Plotting $a_{R}(h)$, Figure 10(c), using Eq. (13), shows values similar to those shown earlier for the geometries with a spherical electrode, but the calculated values of $a_{R}$ are somewhat larger with the needle geometry (in the figure these values have been divided by 2). Since $a_{R}=e\mu/U_{\text{eff}}$, a larger $a_{R}$ would result from a larger mobility or smaller energy cost for producing an electron.
@@ 5 Conclusions
In this work, we considered quasi-static discharge events in air, with a series victim load of relatively small resistance (0.1-0.25 $\Omega$). By measuring both current and voltage profiles as functions of time, we characterized the fraction of stored energy being delivered to the victim load. These measurements were made with several capacitances and gap lengths (which varied the stored energy), two different values of inductance, and several electrode geometries.
@@
With the spherical electrode geometry, varying the capacitance, gap length, and inductance for gap lengths larger than 1 mm produced remarkably consistent results: the fraction of stored energy delivered to the victim load was approximately proportional to both the load resistance and the storage capacitance, and was independent of the gap length (Figure 7). The simple theory Rompe-Weizel model led to $a_{R}$ values between 0.8 - 1.5 cm2/sV2, consistent with what has been reported elsewhere in the literature.
At low gap lengths, there is a significant departure from the scaling described above, in that the ratio of stored energy delivered to the victim load was larger than expected. Seeing that there is an offset in the breakdown voltage at zero gap length, the stored energy does not trend to zero as the gap length goes to zero. We suspect that this breakdown voltage offset may be due to dielectric layers and/or surface texture on the electrodes.
@@
Varying the geometry produces more deviation from simple theory. In particular, the sharp features in the needle-plane/plane-needle geometries significantly decrease the breakdown voltage and energy delivered to the victim. The field concentration at the sharp electrode allows the field to reach breakdown levels at lower potentials. We find that the $a_{R}$ parameter is substantially higher for this geometry. In ongoing work, we see that the channel radius is likely smaller for these ESD events. The deposition of energy into a smaller channel volume likely leads to significantly higher mobility or lower energy cost to generate electrons.
@@
Our results here provide simple scaling for energy delivery: the fraction of the stored energy delivered to the victim is $\eta_{v}<0.5\ (\Omega\mathrm{nF})^{-1}{CR_{v}}$, where $C$ is the capacitance of the system, and $R_{v}$ is the resistance of the victim. As an example, on the high end of the parameters studied here, a 700 pF capacitance and a 0.25 $\Omega$ resistance result in approximately 8% of the total stored energy to be delivered to the victim, independent of gap lengths studied. While the risk that a given absolute energy that is delivered to a sensitive component must be assessed in each situation, we can say that when the resistance of that component is small compared to the resistance of the spark, the vast majority of the stored energy will be deposited into the spark rather than the victim load. Even so, increases in the capacitance and gap distance will generally lead to larger amounts of stored energy, resulting in an increase in the absolute energy delivered to the victim load and, therefore, an increase in the risk to components. In this work, we present investigations of energy transfer in the limit where the victim load resistance is small. In ongoing work, we are investigating the scenarios for arbitrary victim load resistances, where the energy partition can skew towards much higher fractions of stored energy delivered to the victim load. For the small victim loads studied here, the upper bound of energy delivery can provide guidance for energy risk tolerances.

### v2

@@ 4.1 Gap length dependence of breakdown voltage
Since the primary objective of this project is to understand what fraction of the stored energy is delivered to a victim load, we made measurements of the breakdown voltage as a function of the electrode gap for all of the discharge conditions. Figure 7(a) shows the measured voltage just before the start of the discharge as a function of gap separation for the SGS system that had external capacitance values of $C_{x}=100$, 200 and 700 pF (shown in shades of blue) and 1.27 cm diameter brass spherical electrodes, as well as the OAS system where the capacitance was 700 pF (orange circles) and 3.75 cm diameter graphite spherical electrodes. The trends are close to linear in $h$ as expected: Figure 7(b) shows the breakdown field calculated by fitting the $h<2~{\rm mm}$ section of the data to a line. In fitting the data, we find voltage offsets of approximately $V_{0}=1.45$ kV at $h=0$. Similar offsets have previously been observed in the literature [Hogg2015, Meng2019] and likely result from a dielectric layer on the electrodes, such as an oxide layer [Farber2023]. The breakdown fields range from 28–34 kV/cm. The breakdown voltage is expected to vary with electrode geometry and material but can also depend on ambient air pressure and humidity, which was not controlled in our experiment. In principle, all of the data for the SGS chamber in Figure 7(a)should show the same $V_{t}h$; the slight observed variability could result from conditioning of the electrode surfaces during the experiments. The data for the OAS (graphite electrodes) show a slightly lower breakdown threshold than for the SGS (brass electrodes) which we attribute to the electrode material.
@@
Figure 7: Demonstrating the effect of varying capacitance on (a) breakdown voltage, and (b) breakdown field strength for the SGS with $C_{x}$ == 100, 200, and 700 pF, $R_{\text{v}}=0.25$ $\Omega$, and brass 1.27 cm diameter electrodes. The OAS data ($C_{x}$ == 700 pF, $R_{\text{v}}=0.00514$ $\Omega$, 3.75 cm graphite sphere electrodes) is overlaid for comparison with the SGS 700 pF trace to observe the effects of circuit inductance. All data points are averaged over 10 discharge events per gap length and normalized to the victim load size of 0.1 $\Omega$. The legend in (b) applies to both plots.
@@
In Figure 8, we show similar data for the breakdown voltage vs. gap length for various electrode geometries over a broader range of gap lengths. The SGS system was used with $C_{x}=100$ pF and 0.10 $\Omega$ CVR. The legend for the different electrode geometries follows the notation anode-cathode; for example, “needle-plane” indicates a needle-shaped anode and a planar cathode.
@@
Figure 8: Demonstrating the effect of varying electrode geometries on (a) breakdown voltage, and (b) breakdown field strength for the SGS with $C_{x}=100$ pF and $R_{\text{v}}=0.1$ $\Omega$. All data points are averaged over 10 discharge events per gap length. Similar $V_{0}$ offsets were found with a fit on the $V_{\text{th}}$ data with a range between 1.3 and 1.8 kV. The legend in (b) detailing anode-cathode electrode pairings applies to both plots.
@@
For the non-needle electrodes, there is a linear increase in breakdown voltage of approximately 26 kV/cm at smaller gap lengths. The breakdown voltage is lower than the linear trend for larger gap distances. We interpret this as due to the fact that for large gap lengths, the field is less uniform and concentrated near the electrodes. The needle-shaped electrode, either as the anode or cathode, emphasizes this effect, significantly reducing the breakdown voltage; this effect reduces the necessary potentials to initiate discharge events in gaps with strongly non-uniform electric field distributions [Bazelyan1998].
@@ 4.2 Varying circuit capacitance
For a given gap separation, an increase in capacitance increases the initial stored energy and charge. The energy transferred to the victim load $\mathcal{E}_{\text{v}}$ is shown in Figure 9(a). As expected, $\mathcal{E}_{\text{v}}$ increases with $h$ and $C_{x}$. As detailed in Section 3, much of the variation shown can be accounted for in the parameter $\overline{\eta}_{\text{v}}$ that normalizes $\mathcal{E}_{\text{v}}$ to the $\mathcal{E}_{0}$, $R_{\text{v}}$ and $C_{x}$, see Eq. (12). Figure 9(b), shows that for $h>1~{\rm mm}$, $\overline{\eta}_{\text{v}}\approx 0.4~\Omega^{-1}$nF-1 for all of the tested capacitances. The fraction of stored energy transferred to the victim load is quite small: even for $C_{x}=700$ pF, $\eta_{\text{v}}$ is only 2.8%. This confirms the utility of the Rompe-Weizel model. There is, however, a rise in $\overline{\eta}_{\text{v}}$ at lower gaps for the lower stored energy cases. This deviation from a constant results from the voltage offset at $h=0$ giving a stored energy that is larger than expected at small gaps.
The rise in $\overline{\eta}_{\text{v}}$ with smaller gaps seen in Figure 9(b) appears to be related to the observed breakdown voltage offset, $V_{0}$: the stored energy has a component $\mathcal{E}_{0}=\frac{1}{2}CV_{0}^{2}$ that is independent of $h$. When we account for this voltage offset, we can use a different approach to plotting the data. Eq. (11) can be adjusted to replace $E_{\text{th}}$ with $(V_{0}+E_{\text{th}}h)/h=V_{\text{th}}/h$, where $V_{\text{th}}$ is the observed breakdown threshold voltage. Making use of this relation, we can instead plot a calculated $a_{R}(h)$:
$a_{R}=\frac{4\,\eta_{\text{v}}\,h^{2}}{R_{\text{v}}C_{x}V_{\text{th}}^{2}}.$ (13)
Figure 9(c) shows this calculated value of $a_{R}(h)$. The values are fairly constant at larger gap lengths, indicating that the voltage offset is likely responsible for the small-gap divergence in Figure 9(b) for the lower capacitance. Using the SGS, we measured discharges as a function of pressure of synthetic dry air, finding that the value of $a_{R}$ is inversely proportional to the pressure, which is consistent with the expected density dependence of the electron mobility. A variation of air pressure of ${\pm}4\%$ corresponds to a variation of the value of $a_{R}$ by ${\pm}0.075$, small compared to the variation seen in Figure 9(b). The slight decrease at smaller gaps for the larger capacitances indicates that for these cases, the larger discharge current may affect the condition of the electrode surfaces to reduce the voltage offset. The $a_{R}$ values are in a range similar to those observed by Jobava et al. [Jobava2000] that are indicated in the shaded region.
@@ 4.3 Varying circuit inductance
[…]
Figure 9(b) displays the normalized energy transfer to the external capacitor and victim load, calculated using Eq. (12). For both the 700 pF SGS and OAS datasets, the normalized energy remained relatively constant across the gap lengths tested. This agreement in discharge dynamics between the two systems indicates that a tenfold increase in inductance does not significantly impact the energy delivered to the victim load. Since most of the stored energy is dissipated into the spark in the early phase of the discharge, the spark resistance during the latter phase of the discharge is constant. In this case, where the dynamics are in a linear regime, $\eta_{\text{v}}$ would be expected to be independent of the inductance.
Figure 9: Demonstrating the effect of varying capacitance on (a) energy dissipated by $R_{\text{v}}$, (b) $\overline{\eta}_{\text{v}}$, and (c) $a_{R}$ for the SGS with $C_{x}$ == 100, 200, and 700 pF, $R_{\text{v}}=0.25$ $\Omega$, and 1.27 cm diameter brass sphere electrodes. The OAS data ($C_{x}$ == 700 pF, CVR == 0.00514 $\Omega$, 3.75 graphite sphere electrodes) is overlaid for comparison with the SGS 700 pF trace to observe the effects of circuit inductance. All data points are averaged over 10 discharge events per gap length and normalized to the victim load size of 0.1 $\Omega$. The legend in (b) applies to all three plots.
Figure 10: (a) Energy dissipated by $R_{\text{v}}$, (b) $\overline{\eta}_{\text{v}}$, and (c) $a_{R}$ for the SGS with $C_{x}=100$ pF and $R_{\text{v}}=0.1$ $\Omega$, demonstrating the effect of varying electrode geometries. All data points are averaged over 10 discharge events per gap length. The legend in (b) detailing anode-cathode electrode pairings applies to all plots. In (c), the data in blue triangles involving a needle electrode have been divided by a factor of 4; their true scale is labeled on the right side of the plot.
@@ 4.4 Varying electrode geometry
The spherical-electrode geometry for small gaps leads to relatively uniform fields between the electrodes. When we increase the gap length and vary the electrode geometry, the shape and finite size of the electrodes result in non-uniformity of the field, as we saw in the breakdown voltages in Figure 8. The consequence of these effects can be seen in Figure 10(a), which shows the energy delivered to the victim load as a function of gap length for several combinations of electrodes and polarities. The electrodes featuring a needle-shaped electrode and the orientation plane-sphere have lower absolute energy transfer and peak current to the victim. In the plane-sphere geometry, we see lower energy transfer when the cathode is spherical: while this difference is not currently fully understood, it should be noted that in this case the electrodes are made of different materials.
Figure 10(b) is the counterpart to Figure 9(b), showing how the fraction of stored energy transferred to the victim load scaled by $C\,R_{\text{v}}$ varies with gap length. The figure clearly shows that even here, the $\overline{\eta}_{\text{v}}$ scaling results in a reasonably constant value for $h>2~{\rm mm}$, in spite of the different discharge dynamics. Interestingly, while the RW model predicts $\overline{\eta}_{\text{v}}$ to be a constant with no gap length dependence, each geometry shows an initial steep decrease for gaps below 1 mm, followed by a more gradually down-sloping linear relationship at longer gap lengths. The variability in the curve, which differs from a straight line, indicates that more interesting physics is present than the RW model represents, especially at the smaller gap lengths.
Plotting $a_{R}(h)$, Figure 10(c), using Eq. (13), shows values similar to those shown earlier for the geometries with a spherical electrode, but the calculated values of $a_{R}$ are somewhat larger with the needle geometry (in the figure these values have been divided by 4). Since $a_{R}=e\mu/U_{\text{eff}}$, a larger $a_{R}$ would result from a larger mobility or smaller energy cost for producing an electron.
@@ 5 Conclusions
In this work, we considered quasi-static discharge events in air (the voltage was raised in a quasi-static fashion while the gap length was held constant), with a series victim load of relatively small resistance (0.1–0.25 $\Omega$). By measuring both current and voltage profiles as functions of time, we characterized the fraction of stored energy being delivered to the victim load. These measurements were made with several capacitances and gap lengths (which varied the stored energy), two different values of inductance, and several electrode geometries. We test the classic Rompe-Weizel model with our experimental results, finding that the simple model works better than one would think, considering all of the mechanisms that contribute to the empirically-measured $a_{R}$ parameter. The range of $a_{R}$ values we find are consistent with what has been seen in previous work. The variations of the $a_{R}$ parameter shows that deeper investigation is warranted. In ongoing work, we are developing models that more explicitly account for the underlying physics.
@@
With the spherical electrode geometry, varying the capacitance, gap length, and inductance for gap lengths larger than 1 mm produced remarkably consistent results: the fraction of stored energy delivered to the victim load was approximately proportional to both the load resistance and the storage capacitance, and was independent of the gap length (Figure 7). The two spherical electrode types tested were of different radius and material; we did not see a strong dependence of our results connected to the sphere radius, indicating that the independence of the results with gap length is be due to the sphere radius being sufficiently large compared to the gap length. The simple theory Rompe-Weizel model led to $a_{R}$ values between 0.8–1.5 cm2/sV2, consistent with what has been reported elsewhere in the literature.
At low gap lengths, a significant deviation from the established scaling of $\overline{\eta}_{\text{v}}$ was observed, as the ratio of stored energy delivered to the victim load was greater than anticipated. For the smaller capacitances (100–200 pF), this can be explained by an offset in the breakdown voltage at a zero-gap length (see Figure 7(a)), which indicates that the stored energy does not approach zero as the gap length is reduced. We hypothesize that this breakdown voltage offset is caused by the presence of thin dielectric layers and/or surface texture on the electrodes. For the larger capacitance (700 pF), the behavior for $\overline{\eta}_{\text{v}}$ follows the simpler model well. Our conjecture is that repeated discharges at the larger currents associated with the increased stored energy could reduce or remove this oxide layer. The role played by electrode surface conditions on the breakdown voltage merits further investigation, which is beyond the scope of this current paper. Nevertheless, it is worth noting that the deviations from the expected behavior occur under conditions where the initial stored energy (and therefore the energy transferred to the victim load) is the lowest.
@@
Varying the geometry produces more deviation from simple theory. In particular, the sharp features in the needle-plane/plane-needle geometries significantly decrease the breakdown voltage and energy delivered to the victim. The field concentration at the sharp electrode allows the field to reach breakdown levels at lower potentials. We find that the $a_{R}$ parameter is substantially higher for this geometry. Preliminary imaging studies in our lab [SchramaCAM2023IoED] indicate that the channel radius is likely smaller for these ESD events. The deposition of energy into a smaller channel volume likely leads to significantly higher mobility or lower energy cost to generate electrons.
@@
Our results here provide simple scaling for energy delivery: the fraction of the stored energy delivered to the victim is $\eta_{\text{v}}<0.5\ (\Omega\mathrm{nF})^{-1}{CR_{\text{v}}}$, where $C$ is the capacitance of the system, and $R_{\text{v}}$ is the resistance of the victim. As an example, on the high end of the parameters studied here, a 700 pF capacitance and a 0.25 $\Omega$ resistance result in approximately 8% of the total stored energy to be delivered to the victim load, independent of gap lengths studied. While the risk that a given absolute energy that is delivered to a sensitive component must be assessed in each situation, we can say that when the resistance of that component is small compared to the resistance of the spark, the vast majority of the stored energy will be deposited into the spark rather than the victim load. Even so, increases in the capacitance and gap distance will generally lead to larger amounts of stored energy, resulting in an increase in the absolute energy delivered to the victim load and, therefore, an increase in the risk to components. In this work, we present investigations of energy transfer in the limit where the victim load resistance is small. In a related paper [schrama2025variableRv], we have investigated the scenarios for arbitrary victim load resistances, where the energy partition can skew toward much higher fractions of stored energy delivered to the victim load. For the small victim loads studied here, the upper bound of energy delivery can provide guidance for energy risk tolerances.